# Perceptual compression

### v1

@@ Appendix B Additional experimental results
Kodak MS-COCO 30k
Refer to caption Refer to caption
Figure 9: PSNR on Kodak and MS-COCO30k
@@ B.2 Qualitative results
In this ablation study, we conduct a detailed analysis to comprehend the effects of employing text as a global embedding. In Figure 12, we present qualitative examples across two bitrates, where we manipulate the conditioning text. At lower bitrates for the spatial conditioning (first and second rows), we observe that textual conditioning has a significant influence on the reconstructed image. It can even alter the semantics of the scene, such as transforming a lighthouse into a castle (second row, third column). Furthermore, we note the positive impact of text conditioning, as image quality substantially deteriorates when text is not used (first row, second column). Conversely, when the bitrate of the local encoding is higher (last row), we notice that text has minimal influence on the reconstructed image.
Reconstruction at different bitrates.
@@
In Figure 13, we show reconstructions of two MS-COCO 30k images using PerCo trained for different targeted spatial rates, ranging from 0.0046 to 0.1279 bpps. The image quality is quite stable across bitrates, while the similarity with the original images increases with the bitrate.
Qualitative analysis of LDM autoencoder reconstruction
@@
When computing the averaged similarity scores on Kodak dataset for the LDM autoencoder, we obtain 0.0775 LPIPS, 0.9131 MS-SSIM and 25.6003 PSNR which is lower than MS-ILLM at 0.1535 bpp for all the scores and lower than MS-ILLM at 0.0806 bpp on PSNR. In Figure 10 we show a Kodak image illustrating that. Even if the LDM autoencoder produces highly detailed images, we observe that reconstruction fidelity is somewhat worse than MS-ILLM at higher bitrates. Indeed, when looking at image details such as the waves, we notice that MS-ILLM exhibits some blurring at 0.0806 bpp while LDM produces sharp details for a similar PSNR. When comparing with MS-ILLM at the higher bitrate of 0.1535 bpp, we notice that the letters on the boat are less faithful for our LDM Autoencoder than MS-ILLM with respect to the original image.
Qualitative analysis of pixelwise similarity metrics
@@
We observed that pixelwise distortion metrics such as PSNR, MS-SSIM and LPIPS are not very meaningful at very low bitrates, as already pointed out by Lei et al. (2023). In the example in Figure 14, PerCo at 0.0031 bpp is 3.33 points worse in PSNR, 0.1766 points worse in MS-SSIM and 0.085 points worse in LPIPS than MS-ILLM at 0.0065 bpp. However, at these very low bit-rates PerCo yields a much sharper and realistic reconstruction than MS-ILLM, and preserves most of the semantics of the scene. These examples provide some indication that these metrics do not align well with human preference for image quality at very low bitrates.
Additional comparisons to baselines
@@
In Figure 15, we compare PerCo at 0.1264 bpp, the highest bitrate we trained on, to VTM, HiFiC and MS-ILLM at similar bitrates. At this bitrate, PerCo faithfully reconstructs the original images with similarly detailed textures as HiFiC and MS-ILLM, and fewer artifacts than VTM.

### v2

@@ Appendix B Additional experimental results
Figure 9: PSNR on Kodak (left) and MS-COCO30 (right).
@@ B.1 Quantitative results
In Figure 10, we provide evaluation on an additional global image similarity metric that consists in computing CLIP scores with a cosine similarity between CLIP image embeddings of original and reconstructed images, instead of cosine similarity between text captions and reconstructed images. This is motivated by the fact that we noticed that CLIP text score was saturating at high bitrates, as it reaches the upper bound of 30.4 when computed on real images instead of reconstructed ones, while CLIP image score threshold is 1 when evaluated on real images which is higher than what is obtained at high bitrates for all the baselines.
Figure 10: Evaluation in terms of the CLIP image metric.
@@ B.2 Qualitative results
Ablations on text. In this ablation study, we conduct a detailed analysis to comprehend the effects of employing text as a global embedding. In Figure 13, we present qualitative examples across two bitrates, where we manipulate the conditioning text. At lower bitrates for the spatial conditioning (first and second rows), we observe that textual conditioning has a significant influence on the reconstructed image. It can even alter the semantics of the scene, such as transforming a lighthouse into a castle (second row, third column). Furthermore, we note the positive impact of text conditioning, as image quality substantially deteriorates when text is not used (first row, second column). Conversely, when the bitrate of the local encoding is higher (last row), we notice that text has minimal influence on the reconstructed image.
@@
Reconstruction at different bitrates. In Figure 14, we show reconstructions of two MS-COCO 30k images using PerCo trained for different targeted spatial rates, ranging from 0.0046 to 0.1279 bpps. The image quality is quite stable across bitrates, while the similarity with the original images increases with the bitrate.
@@
Qualitative analysis of LDM autoencoder reconstruction. When computing the averaged similarity scores on Kodak dataset for the LDM autoencoder, we obtain 0.0775 LPIPS, 0.9131 MS-SSIM and 25.6003 PSNR which is lower than MS-ILLM at 0.1535 bpp for all the scores and lower than MS-ILLM at 0.0806 bpp on PSNR. In Figure 11 we show a Kodak image illustrating that. Even if the LDM autoencoder produces highly detailed images, we observe that reconstruction fidelity is somewhat worse than MS-ILLM at higher bitrates. Indeed, when looking at image details such as the waves, we notice that MS-ILLM exhibits some blurring at 0.0806 bpp while LDM produces sharp details for a similar PSNR. When comparing with MS-ILLM at the higher bitrate of 0.1535 bpp, we notice that the letters on the boat are less faithful for our LDM Autoencoder than MS-ILLM with respect to the original image.
@@
Qualitative analysis of pixelwise similarity metrics. We observed that pixelwise distortion metrics such as PSNR, MS-SSIM and LPIPS are not very meaningful at very low bitrates, as already pointed out by Lei et al. (2023). In the example in Figure 15, PerCo at 0.0031 bpp is 3.33 points worse in PSNR, 0.1766 points worse in MS-SSIM and 0.085 points worse in LPIPS than MS-ILLM at 0.0065 bpp. However, at these very low bit-rates PerCo yields a much sharper and realistic reconstruction than MS-ILLM, and preserves most of the semantics of the scene. These examples provide some indication that these metrics do not align well with human preference for image quality at very low bitrates.
@@
Additional comparisons to baselines. In Figure 16, we compare PerCo at 0.1264 bpp, the highest bitrate we trained on, to VTM, HiFiC and MS-ILLM at similar bitrates. At this bitrate, PerCo faithfully reconstructs the original images with similarly detailed textures as HiFiC and MS-ILLM, and fewer artifacts than VTM.